# Argument Rarity-based Originality Assessment for AI-Assisted Writing

### v1

@@ 4.2 Evaluation on Originality Score Distribution
This section analyzes the originality score distributions of human essays and AI essays to verify whether the proposed framework AROA can appropriately evaluate originality.
@@ 4.2.1 Overall Score Distribution
Figure 2 shows the distribution of final originality scores $S_{\text{final}}$ for human essays ($n=1,375$) and AI essays ($n=1,000$). Both distributions overlap around the mean, with the human essay mean at 0.015-0.015 and the AI essay mean at 0.001-0.001, which are nearly equivalent. What is noteworthy is the difference in distribution shapes. Human essays are concentrated in a relatively narrow range (standard deviation 0.345), while AI essays are dispersed over a wider range (standard deviation 0.534). This suggests that AI essay originality varies greatly depending on the prompting and sampling conditions.
@@
Table 3 summarizes the statistical comparison. The final originality score $S_{\text{final}}$ showed no statistically significant difference between human essays and AI essays ($p=0.118$, Cohen’s d=0.032d=-0.032). This result has important implications. The proposed framework does not function as a mere AI detector but evaluates the originality of argumentation itself regardless of author attribution. Even human-written essays receive low scores if they remain at typical argumentation, while AI-generated essays receive high scores if they contain original argumentation.
@@ 4.2.2 Component-wise Analysis
First, in claim rarity $S_{\text{claim}}$, human essays showed substantially higher scores (human: $0.170$, AI: $0.037$, $d=2.24$). As shown in the second panel of Figure 3, this difference is visually striking, with AI essay scores remaining at approximately one-fifth of human essays. This indicates higher diversity of perspectives and claims in human essays. AI-generated essays tend to converge on typical claim patterns contained in training data, resulting in a clustered distribution in semantic space.
A similar trend was observed in evidence rarity $S_{\text{evid}}$ (human: $0.092$, AI: $0.048$, $d=1.34$). Human essays used more diverse examples, statistics, and citations, while AI essays relied on commonly cited typical evidence. As the error bars indicate, human essays also show greater variation in evidence selection, suggesting the inclusion of unique evidence based on individual knowledge and experience.
On the other hand, no substantial differences were observed in structural rarity $S_{\text{struct}}$ (human: $4.987$, AI: $4.797$, $d=0.20$) and cognitive depth $S_{\text{move}}$ (human: $5.093$, AI: $5.156$, d=0.03d=-0.03). As shown in the first and fourth panels of Figure 3, the bar graphs for both groups are nearly the same height, and the error bars overlap. This indicates that AI can sufficiently reproduce the structural complexity in human argumentation (hierarchical structure of claim-reason-evidence and dialogic structure of counterargument-rebuttal).
@@ 4.3 Evaluation on Quality-Originality Relationship
This section analyzes the relationship between quality and originality and verifies the validity of the multi-component design of the proposed framework AROA.
@@ 4.3.1 Component Independence
Figure 4 shows the correlation matrix among originality components and quality score. First, structural rarity $S_{\text{struct}}$ has extremely low correlations with all other components (r=0.08r=0.080.130.13). This indicates that the atypicality of argument structure is a dimension independent of claim/evidence content and depth of reasoning. In other words, essays with structurally rare argumentation patterns do not necessarily contain rare claims or evidence. This result supports the design philosophy of this framework that how arguments are made and what is argued are different evaluation axes.
Next, cognitive depth $S_{\text{move}}$ also has low correlations with other components (r=0.06r=0.060.130.13). The dialogic structure of counterargument-rebuttal and hierarchical depth of reasoning vary independently of claim/evidence rarity. Essays with deep reasoning structures do not necessarily make rare claims, and vice versa.
On the other hand, a moderate positive correlation ($r=0.51$) was observed between claim rarity $S_{\text{claim}}$ and evidence rarity $S_{\text{evid}}$. This is an intuitively understandable result. Authors who make claims from original perspectives tend to also explore and present unique evidence to support those claims. Both components share the common property of semantic rarity while evaluating different argument elements (claims vs. evidence), so they do not completely coincide.
These results support the validity of multifaceted evaluation using four components. If all components showed high correlations, the significance of using multiple components would diminish. However, in practice, each component captures different aspects of originality, enabling multifaceted evaluation that is not visible with a single indicator.
@@ 4.3.2 Quality Does Not Guarantee Originality
Behind this negative correlation lies a trade-off between quality and rarity. To write logically coherent, high-quality text, it is easier to use widely accepted claim patterns and commonly cited evidence. Using original claims and rare evidence makes it difficult to maintain logical coherence, and quality scores tend to decrease.
@@
These results clearly demonstrate that text quality does not guarantee originality. AI essays generate fluent and logically coherent text, but their high quality is achieved by relying on typical claim patterns and common evidence. The strong negative correlation between $Q$ and $S_{\text{claim}}$$S_{\text{evid}}$ shown in the correlation matrix quantitatively supports this trade-off. The reason AI essays that would receive high scores in conventional writing assessment remain at originality scores equivalent to human essays in this framework is that it appropriately captures this quality-rarity trade-off. The proposed framework appropriately models the relationship between quality and originality through the design that multiplies quality score $Q$ into the final score (Equation 18). In AI essays, the low originality components are partially compensated by high $Q$, while in human essays, high originality components are adjusted by the variation in $Q$. As a result, the final score distributions of both groups become similar, reflecting that quality and originality are independent evaluation axes.
@@ 4.3.3 Illustrative Examples
There are three reasons why this essay received a high score. First, it presents original perspectives such as stress reduction and increased exercise, not just environmental issues, leading to a high claim rarity of $S_{\text{claim}}=0.294$. Second, by raising counterarguments (weather, long-distance commuting, transporting work tools) and responding to them, it achieves a deep cognitive depth of $S_{\text{move}}=10.0$. Third, although the quality score is not perfect at $Q=0.75$, the original content compensates for it.
@@
This essay is logically coherent and achieves $Q=1.00$, but has the lowest originality score. There are three reasons for this. First, claims such as “good for the environment” and “good for the community” are stereotypical patterns frequently seen in other essays addressing the same topic, leading to a low claim rarity of $S_{\text{claim}}=0.038$. Second, throughout the essay, it takes the form of listing citations from materials, lacking original analysis or specific examples ($S_{\text{evid}}=0.046$). Third, it does not consider any counterarguments to its claims and remains at a simple “reason 1, reason 2” structure, resulting in the minimum cognitive depth of $S_{\text{move}}=3.0$.
These concrete examples clearly show what the proposed framework evaluates. High originality scores arise from original perspectives, diverse evidence, and deep argumentation structures that anticipate and respond to counterarguments. On the other hand, essays that are logically coherent but rely on typical claims and common evidence receive low scores. This succinctly demonstrates the difference between conventional writing assessment (quality-focused) and this framework (originality-focused).
@@ 4.4 Evaluation on Robustness to Sample Size
This section verifies the stability of scores against reference corpus size. Since the proposed method is based on density estimation, the reference corpus size may affect score reliability. Using the entire human essay dataset as a reference, subsampling experiments were conducted. For each sample size $n\in\{30,50,100,200,300,500,700,1000\}$, the following procedure was repeated 5 times with different random seeds: i) randomly sample $n$ essays from the full dataset, ii) recalculate structural rarity scores $S_{\text{struct}}$ using only the subsample, and iii) compute Spearman’s rank correlation coefficient $\rho$ between rankings based on the subsample and rankings based on the full data. $S_{\text{struct}}$, which depends on density estimation, is the component most sensitive to sample size, and if this is stable, other components can be expected to be similarly stable.
@@ 4.5 Cost Analysis
This section analyzes computational cost and API cost for practical deployment of the proposed framework. The experiments were conducted on a macOS environment equipped with an Apple M1 chip (8-core CPU, 7-core GPU) and 16GB RAM, using 10 parallel workers.
Table 5 shows the processing time breakdown per essay. Approximately 90% of processing time is occupied by five-element extraction through LLM API calls. This reflects that the process of analyzing the entire essay and structuring claims, reasons, evidence, counterarguments, and rebuttals is the most computationally intensive. On the other hand, text embedding and density estimation are negligible as local processing, making it an efficient design from an API cost perspective as well.
@@
In comparison with human evaluation, the efficiency of the proposed method is remarkable. When human evaluators manually evaluate essay originality, it typically takes 5–10 minutes per essay. In contrast, the proposed method can process in approximately 1.5 seconds, which is 200–400 times faster. In terms of cost, approximately \$0.0024 per essay compares to \$1–5 for human evaluation depending on evaluator expertise and evaluation depth, making it approximately 400–2,000 times cheaper. This efficiency enables large-scale and frequent originality assessment that was previously difficult to implement.
@@ 4.6 Evaluation on Transferability Across LLMs
This section conducts transferability analysis using three models, GPT-4.1-mini, Gemini 2.5 Flash Lite, and Claude 3 Haiku, to verify whether the proposed framework depends on the specific LLM used for argument structure extraction.
Table 7 shows the cross-model correlations of final originality score $S_{\text{final}}$. The average Pearson correlation across all model pairs was $r=0.465$, confirming moderate agreement in originality evaluation. GPT showed relatively high agreement with both Gemini and Claude, while the correlation between Gemini and Claude was the lowest ($r=0.402$), suggesting that these two models emphasize different aspects of argumentative originality.
@@
Table 8 shows the cross-model agreement for each component score. Notable differences in transferability were observed depending on the component. Evidence rarity $S_{\text{evid}}$ showed the highest stability (average $r=0.616$), confirming that evidence identification and evaluation are relatively consistent across different LLMs. On the other hand, structural rarity $S_{\text{struct}}$ showed the lowest agreement (average $r=0.232$), revealing that argument structure recognition differs significantly by LLM. Quality evaluation $Q$ also showed low transferability (average $r=0.249$). In particular, Claude assigned nearly uniform quality scores to almost all essays, suggesting that quality evaluation criteria fundamentally differ between models.
@@
Figure 7 shows the contribution (correlation coefficient) of each component to $S_{\text{final}}$ for each model. Characteristic evaluation tendencies were observed for each model. GPT-4.1-mini shows nearly equal contributions across all four components (r=0.54r=0.540.560.56), performing balanced evaluation. Gemini 2.5 Flash Lite emphasizes structural rarity $S_{\text{struct}}$ the most ($r=0.55$) while showing relatively low sensitivity to evidence rarity $S_{\text{evid}}$ ($r=0.43$). Claude 3 Haiku places the greatest weight on claim rarity $S_{\text{claim}}$ ($r=0.62$) and evidence rarity $S_{\text{evid}}$ ($r=0.58$), showing a pronounced tendency to emphasize semantic rarity.
@@
These results indicate that the proposed framework captures the core constructs of argumentative originality regardless of the LLM adopted, while revealing that component stability differs systematically. Evidence-related scores transfer well across models, while structure and quality evaluations are more model-dependent. The differences in evaluation characteristics of each model and their practical implications are discussed in Section 5.
@@ 5.1 Key Findings
This study proposed AROA, a framework for automatically evaluating argumentative originality in essays, and verified its effectiveness through experiments using 1,375 human essays and 1,000 AI essays. This section discusses the key findings from the experimental results and their implications.
The most important finding is that text quality and originality are independent evaluation axes, and a trade-off exists between them. In the correlation analysis, strong-to-moderate negative correlations were observed: r=0.67r=-0.67 between quality score $Q$ and claim rarity $S_{\text{claim}}$, and r=0.50r=-0.50 between $Q$ and evidence rarity $S_{\text{evid}}$. This result suggests that to write logically coherent, high-quality text, it is easier to use widely accepted claim patterns and common evidence. Using original claims and rare evidence makes justification and maintenance of logical coherence difficult, resulting in lower quality scores. This finding highlights the problem of conventional writing assessment being biased toward quality. Fluent and logically coherent text does not necessarily reflect original thinking; rather, it may indicate dependence on typical patterns.
The comparison between human essays and AI essays vividly demonstrates this quality-originality trade-off. While AI essays achieved near-perfect quality scores ($Q=0.998$), their claim rarity remained at approximately one-fifth of human essays (AI: 0.037, human: 0.170). Interestingly, no significant differences were observed between the two groups in structural rarity $S_{\text{struct}}$ and cognitive depth $S_{\text{move}}$. This means that contemporary LLMs can sufficiently reproduce human argumentation structures—the hierarchical structure of claim-reason-evidence and the dialogic structure of counterargument-rebuttal. However, in the semantic dimension of what to claim and what evidence to choose, AI tends to converge on typical patterns contained in training data, finding it difficult to generate unique evidence based on diverse perspectives and personal experiences that humans possess. This finding provides a perspective different from AI-generated text detection. The proposed framework AROA does not determine whether the author is AI or human but evaluates the originality of argumentation itself. In fact, no statistically significant difference was observed in the final score $S_{\text{final}}$ between the two groups ($p=0.118$). This means that even human-written essays receive low scores if they remain at typical argumentation, while AI-generated essays receive high scores if they contain original argumentation.
The validity of multifaceted evaluation using four components was also confirmed. In the correlation analysis, structural rarity $S_{\text{struct}}$ and cognitive depth $S_{\text{move}}$ showed extremely low correlations with other components (r=0.06r=0.060.130.13), indicating that each captures different aspects of originality. How arguments are made (structure) and what is argued (semantic content) are independent evaluation axes, enabling multifaceted originality evaluation that cannot be captured by a single indicator. On the other hand, a moderate positive correlation ($r=0.51$) was observed between claim rarity $S_{\text{claim}}$ and evidence rarity $S_{\text{evid}}$. This indicates that authors who make claims from original perspectives tend to also explore and present unique evidence to support those claims, confirming that while sharing the common property of semantic rarity, they evaluate different argument elements.
@@
From a practical perspective, this framework has sufficient efficiency and reliability for deployment in educational settings. Processing costs of approximately 1.5 seconds and approximately \$0.0024 per essay make large-scale evaluation realistic. Robustness analysis showed that rank correlation $\rho\geq 0.80$ is achieved when the reference corpus size is 300 or more, yielding reliable results for undergraduate-level courses and evaluations pooling multiple classes. This finding provides a rationale for large-scale course designs. The RaaS implementation at Ritsumeikan University, which accommodates up to 1,000 students per class as mentioned in Section 1, is well-suited for density-based originality assessment. Larger cohorts provide richer reference corpora, enabling more reliable rarity estimation. In this sense, the seemingly challenging scale of large enrollment courses becomes an advantage rather than a limitation when combined with AI-assisted originality assessment. On the other hand, at typical classroom scales (30–50 essays), rank correlation remains at approximately 0.50, so for small-scale classes, it is recommended to use the framework for formative feedback rather than summative ranking.
@@
Transferability analysis confirmed that the proposed framework does not heavily depend on a specific LLM. Among the three models—GPT-4.1-mini, Gemini 2.5 Flash Lite, and Claude 3 Haiku—the average correlation of final scores was $r=0.465$, showing moderate agreement. By component, evidence rarity $S_{\text{evid}}$ showed the highest stability (average $r=0.616$), while structural rarity $S_{\text{struct}}$ and quality evaluation $Q$ showed high model dependency (average r=0.23r=0.230.250.25). Each model showed characteristic evaluation tendencies: GPT performed balanced evaluation, Gemini emphasized structure, and Claude emphasized semantic rarity. This result suggests the possibility of more robust evaluation through an ensemble of multiple models.
These findings suggest a paradigm shift in writing assessment in the AI era. Now that LLMs can easily generate high-quality text, assessment based solely on quality is losing its significance. What is important is whether the author has original perspectives, explores diverse evidence, and has gone through a deep thinking process that anticipates and responds to counterarguments. However, it is not easy for human evaluators to consistently evaluate such originality. When grading numerous essays, evaluators are susceptible to order effects and fatigue, where impressions from earlier essays influence subsequent evaluations. Also, evaluators tend to give higher scores to claims that match their prior knowledge and expectations, while finding it difficult to appropriately recognize the value of unfamiliar perspectives or rare evidence. Furthermore, considering relative rarity within the entire corpus when evaluating individual essays is cognitively difficult. The proposed framework AROA complements these limitations of human evaluation. Through density-based rarity evaluation, it can objectively and consistently quantify how rare each essay’s claims and evidence are within the entire corpus. It can evaluate all essays according to the same criteria without being affected by evaluator fatigue or order effects. This provides a foundation for educators to more fairly and efficiently evaluate and develop students’ critical thinking skills while referring to the proposed framework’s scores.
5.2 Limitations
This study has several limitations. First, the evaluation is based on a single dataset and is limited to two topics. Generalizability to different topics, different languages, and different grade levels requires future verification. In particular, applicability to genres other than argumentative essays, such as specialized academic fields and creative writing, remains unverified.
Second, the proposed framework has not been validated for correlation with originality evaluation by human experts. This study theoretically discussed that density-based rarity is a valid operational definition of originality and demonstrated its validity through concrete examples, but future empirical research is needed to determine how well this framework’s scores align with essays that educators and researchers intuitively judge as original.
Third, since argument structure extraction and quality evaluation depend on LLMs, LLM biases and limitations may affect scores. Although transferability analysis confirmed moderate cross-model agreement, high model dependency was shown particularly in structural rarity and quality evaluation. Also, since LLM capabilities are rapidly evolving, the results from models used in this study may not directly apply to future models.
Fourth, there is dependency on reference corpus size. Robustness analysis showed that reliable evaluation is possible with 300 or more essays, but there are constraints on application in small-scale classes. Further consideration is also needed regarding how the composition of the reference corpus (students’ grade levels, assignment difficulty, etc.) may affect scores.

### v3

@@ 4.2.1 Overall Score Distribution
Figure 2 shows the distribution of final originality scores $S_{\text{final}}$ for human essays ($n=1,375$) and AI essays ($n=1,000$). Both distributions overlap around the mean (human: 0.015-0.015, AI: 0.001-0.001), but differ in shape: human essays are concentrated in a narrow range (SD = 0.345), while AI essays are more dispersed (SD = 0.534), suggesting that AI originality varies with prompting and sampling conditions.
@@
Table 3 summarizes the statistical comparison. The final originality score $S_{\text{final}}$ showed no statistically significant difference between human essays and AI essays ($p=0.118$, Cohen’s d=0.032d=-0.032). This indicates that AROA evaluates the originality of argumentation itself rather than functioning as an AI detector: human essays receive low scores if argumentation is typical, and AI essays receive high scores if argumentation is original.
@@ 4.2.2 Component-wise Analysis
In claim rarity $S_{\text{claim}}$, human essays scored substantially higher (human: $0.170$, AI: $0.037$, $d=2.24$), with AI scores at approximately one-fifth of human levels. AI-generated essays tend to converge on typical claim patterns from training data, resulting in a clustered distribution in semantic space.
A similar trend was observed in evidence rarity $S_{\text{evid}}$ (human: $0.092$, AI: $0.048$, $d=1.34$), with human essays showing more diverse evidence selection while AI essays relied on commonly cited evidence.
In contrast, no substantial differences were observed in structural rarity $S_{\text{struct}}$ ($d=0.20$) and cognitive depth $S_{\text{move}}$ (d=0.03d=-0.03), indicating that AI can reproduce the structural complexity of human argumentation.
@@ 4.3.1 Component Independence
Figure 4 shows the correlation matrix among originality components and quality score. First, structural rarity $S_{\text{struct}}$ has extremely low correlations with all other components (r=0.08r=0.080.130.13). This indicates that structural atypicality is independent of claim/evidence content and depth of reasoning, supporting the design philosophy that how arguments are made and what is argued are different evaluation axes.
Cognitive depth $S_{\text{move}}$ also has low correlations with other components (r=0.06r=0.060.130.13), confirming that reasoning depth varies independently of semantic rarity.
A moderate positive correlation ($r=0.51$) between $S_{\text{claim}}$ and $S_{\text{evid}}$ suggests that authors with original claims tend to also present unique evidence, though the components evaluate different argument elements and do not fully coincide.
These results support the validity of multifaceted evaluation, as each component captures different aspects of originality.
@@ 4.3.2 Quality Does Not Guarantee Originality
This reflects a quality-rarity trade-off: logically coherent, high-quality text is easier to produce using widely accepted claim patterns, while original claims and rare evidence make maintaining coherence more difficult.
@@
These results demonstrate that text quality does not guarantee originality. AI essays achieve high quality by relying on typical claim patterns, while human essays show higher semantic originality but lower quality scores. The multiplicative design (Equation 18) ensures that low originality in AI essays is partially compensated by high $Q$, while high originality in human essays is adjusted by variation in $Q$, resulting in similar final score distributions and confirming that quality and originality are independent axes.
@@ 4.3.3 Illustrative Examples
This essay received a high score because it presents original perspectives beyond environmental issues ($S_{\text{claim}}=0.294$), raises and responds to counterarguments ($S_{\text{move}}=10.0$), and its original content compensates for imperfect quality ($Q=0.75$).
@@
Despite perfect quality ($Q=1.00$), this essay scores low because it relies on stereotypical claims ($S_{\text{claim}}=0.038$), lists citations without original analysis ($S_{\text{evid}}=0.046$), and lacks counterarguments ($S_{\text{move}}=3.0$).
These examples illustrate the distinction between conventional quality-focused assessment and AROA’s originality-focused evaluation.
@@ 4.4 Evaluation on Robustness to Sample Size
Since the proposed method is based on density estimation, the reference corpus size may affect score reliability. Subsampling experiments were conducted for sample sizes $n\in\{30,50,100,200,300,500,700,1000\}$, repeated 5 times each. For each subsample, $S_{\text{struct}}$ (the component most sensitive to sample size) was recalculated and Spearman’s $\rho$ was computed against full-data rankings.
@@ 4.5 Cost Analysis
The experiments were conducted on a macOS environment equipped with an Apple M1 chip (8-core CPU, 7-core GPU) and 16GB RAM, using 10 parallel workers. Table 5 shows the processing time breakdown per essay. Approximately 90% of processing time is occupied by argument extraction through LLM API calls, while embedding and density estimation are negligible as local processing.
@@
Compared to human evaluation (5–10 minutes and \$1–5 per essay), the proposed method is 200–400 times faster (1.5 seconds) and 400–2,000 times cheaper (\$0.0024), enabling large-scale originality assessment that was previously impractical.
@@ 4.6 Evaluation on Transferability Across LLMs
To verify whether the proposed framework depends on the specific LLM used, transferability analysis was conducted using GPT-4.1-mini, Gemini 2.5 Flash Lite, and Claude 3 Haiku.
The average cross-model Pearson correlation for $S_{\text{final}}$ was $r=0.465$ (Table 7), confirming moderate agreement. GPT showed relatively high agreement with both Gemini and Claude, while Gemini–Claude correlation was lowest ($r=0.402$).
@@
Component-wise analysis (Table 8) revealed that evidence rarity $S_{\text{evid}}$ showed the highest stability (average $r=0.616$), while structural rarity $S_{\text{struct}}$ (average $r=0.232$) and quality $Q$ (average $r=0.249$) showed low transferability. Claude in particular assigned nearly uniform quality scores, suggesting fundamentally different evaluation criteria.
@@
Each model showed characteristic evaluation tendencies (Figure 7): GPT contributed equally across components (r=0.54r=0.540.560.56), Gemini emphasized structural rarity ($r=0.55$), and Claude emphasized semantic rarity ($S_{\text{claim}}$: $r=0.62$, $S_{\text{evid}}$: $r=0.58$).
@@
These results indicate that AROA captures argumentative originality regardless of the LLM adopted, though component stability varies: evidence scores transfer well, while structure and quality evaluations are more model-dependent.
@@ 5.1 Key Findings
The experiments using 1,375 human essays and 1,000 AI essays yielded three key findings. First, a strong negative correlation between text quality and semantic rarity (r=0.67r=-0.67 for $Q$ vs. $S_{\text{claim}}$; r=0.50r=-0.50 for $Q$ vs. $S_{\text{evid}}$) reveals a quality-originality trade-off: logically coherent, high-quality text is easier to produce using widely accepted claims and common evidence, while original claims make maintaining coherence more difficult.
Second, human essays and AI essays exhibit markedly different component profiles despite similar final scores ($p=0.118$). AI essays achieved near-perfect quality ($Q=0.998$) but their claim rarity remained at approximately one-fifth of human levels (AI: 0.037, human: 0.170). In contrast, no significant differences were observed in structural rarity ($d=0.20$) or cognitive depth (d=0.03d=-0.03). This indicates that contemporary LLMs can reproduce the structural complexity of human argumentation but tend to converge on typical semantic patterns from training data.
Third, the four originality components capture genuinely different aspects of originality. Structural rarity and cognitive depth showed low correlations with other components (r=0.06r=0.060.130.13), while claim and evidence rarity showed a moderate correlation ($r=0.51$), confirming that how arguments are made and what is argued are independent evaluation axes.
5.2 Theoretical Implications
These findings contribute to the theoretical understanding of writing assessment and AI-generated text in several ways. First, the quality-originality trade-off challenges the longstanding assumption in AES research that text quality is the primary evaluation target. Our results demonstrate that quality and originality are not only independent but inversely related in their semantic dimensions, suggesting that conventional quality-focused assessment systematically undervalues original thinking. This extends the creativity assessment literature [Runco and Yoruk, 2014, Beaty et al., 2022], which has established rarity as a measure of originality for short responses in divergent thinking tests, to the domain of complex argumentative writing.
Second, the component analysis provides empirical evidence for the multidimensional nature of argumentative originality. The low correlations between structural and semantic components support the theoretical distinction between the form and content of argumentation, consistent with Toulmin’s framework [Toulmin, 2003]. This finding implies that originality assessment frameworks must adopt multifaceted approaches rather than relying on single indicators.
Third, the human-AI comparison reveals that the limitations of current LLMs lie not in structural competence but in semantic originality. This finding contributes to the growing literature on AI capabilities and limitations by identifying a specific dimension, content originality, where human writers retain a clear advantage. Unlike AI detection approaches that focus on surface-level artifacts, AROA provides a content-level framework for understanding the qualitative differences between human and AI writing.
5.3 Practical Implications
@@
From a practical perspective, AROA has sufficient efficiency for deployment in educational settings, with processing costs of approximately 1.5 seconds and \$0.0024 per essay. Robustness analysis showed that rank correlation $\rho\geq 0.80$ is achieved with a reference corpus of 300 or more essays, yielding reliable results for courses pooling multiple classes. The RaaS implementation at a large Japanese university, which accommodates up to 1,000 students per class [Ritsumeikan University Institute for Teaching and Learning, 2025], exemplifies settings well-suited for density-based originality assessment: larger cohorts provide richer reference corpora, making the scale of large enrollment courses an advantage rather than a limitation. For small-scale classes (30–50 essays), where rank correlation remains at approximately 0.50, the framework is recommended for formative feedback rather than summative ranking.
@@
Transferability analysis confirmed that AROA does not heavily depend on a specific LLM ($r=0.465$ average cross-model correlation for $S_{\text{final}}$). Evidence rarity showed the highest stability ($r=0.616$), while structural rarity and quality evaluation were more model-dependent (r=0.23r=0.230.250.25). Each model showed characteristic evaluation tendencies: GPT performed balanced evaluation, Gemini emphasized structure, and Claude emphasized semantic rarity, suggesting the possibility of more robust evaluation through model ensembles.
These results suggest that AROA can complement the limitations of human evaluation in assessing originality. Human evaluators face order effects and fatigue [Mahshanian and Shahnazari, 2020], expertise bias [Alvén, 2019], and cognitive limitations in judging relative rarity across entire corpora. AROA addresses these by providing objective, consistent, corpus-wide rarity quantification, offering educators a foundation for more fairly evaluating and developing students’ critical thinking skills.
5.4 Limitations and Future Work
This study has several limitations. First, the evaluation is based on a single dataset limited to two topics, and generalizability to different topics, languages, grade levels, and genres (e.g., academic writing, creative writing) requires future verification. Second, the framework has not been validated against originality evaluations by human experts; empirical research is needed to determine how well AROA scores align with expert judgments of originality. Third, since argument extraction and quality evaluation depend on LLMs, model biases may affect scores. Although transferability analysis confirmed moderate cross-model agreement, structural rarity and quality evaluation showed high model dependency. Fourth, reliable evaluation requires a reference corpus of 300 or more essays, constraining application in small-scale classes.